Solve each of these systems using the substitution method.
2x + 3y = -4
x - 2y = 5
Any help will be Greatly Appreciated!

Answer by 10301650(32)

GIVEN:
2x+3y=-4
x-2y=5
--(use the second given to look for the X)
x-2y=5
(transfer the -2y to the other side)
x=2y+5 (use this one to solve or the X)
--(now use the first given to substitute the second given)
2x+3y=-4
2(2y+5)+3y=-4
4y+10+3y=-4
7y+10=-4
7y=-10-4
7y=-14
(divide both sides with 7)
y=-2 (answer for Y)
--(NOW, use the second given to solve for X)
x-2y=5
x-2(-2)=5
x+4=5
x=-4+5
x=1(answer for X)